A PARTICLE MOVES ALONG A STRAIGHT LINE WITH UNIFORM ACCELERATION. IF P1 AND P2 ARE LINEAR MOMENTA WHILE CROSSING THE POINTS A AND B, WHAT IS MOMENTAUM WHILE CROSSING C WHICH IS THE MID POINT OF A AND ?????.....
 
can u  give me proof sir

Grade:12th pass

1 Answers

Shailendra Kumar Sharma
6 years ago
Uniform acceleration, let say the acceleration be a, and a will be constant.
Now dv/dt =a 
assume he mass of object m and at point A assume the velocity be v1 and at the point B to velcity be v2
Now distance between A and B can be calculated by 
v2=u2+2as
so v22 =v12 + 2aS
 so the distance between A and B will be S=(v22 – v12)/(2a)
at the midpoint, the distance will be S/2 so C is s/2 far from A
so velocity at C

vc2 =v12 + 2aS = v12  + 2a{(v22 – v12)/(2a)}(/2)
=v12 +(v22 – v12)/2=(v22 + v12)/2
So vc2 = (v22 + v12)/2 multiple by m2 both side
m2 vc2 = m2 (v22 + v12)/2
Pc2 = (PA2 +PB2)/2